OpenStudy (anonymous):

PLEASE HELP URGENT!! Chase wants to factor x^2 + 10x + 25 by grouping; however, Paige says it is a special product and can factor a different way. Using complete sentences, explain and demonstrate how both methods will result in the same factors.

OpenStudy (welshfella):

factor by grouping: x^2 + 10x + 25 = x^2+ 5x + 5x + 25 = x(x + 5) + 5 (x + 5) = (x + 5)( x + 5)

OpenStudy (anonymous):

thank you so much for that :)

OpenStudy (anonymous):

what does a special product mean and what is another method I can use ? @welshfella

OpenStudy (welshfella):

well I guess because 25 is a perfect square and we have 2*5 = 10 for the coefficient of x, we can use the following special product identity: (x + a)^2 = x^2 + 2ax + a^2 here a^2 = 25 so a = 5 and 2a = 2*5 = 10

OpenStudy (welshfella):

another example of en expression where we could apply the same identity would b x^2 + 6x + 9

OpenStudy (welshfella):

9 =3^2 and 2*3 = 6

OpenStudy (anonymous):

Oh ok I understand now, thanks a bunch! :)

OpenStudy (welshfella):

yw